OpenStudy (anonymous):

rationalize the denominator and simplify √3:((5a^5)/(3b))

OpenStudy (zzr0ck3r):

im confused

OpenStudy (zzr0ck3r):

can you draw the expression

OpenStudy (zzr0ck3r):

\[\sqrt{3}:\frac{5a^5}{3b}\]?

OpenStudy (anonymous):

\[\sqrt[3]{\frac{ 5a^5 }{ 3b }}\] This is the question rationalize this denominator and simplify

OpenStudy (anonymous):

@zzr0ck3r yes

OpenStudy (anonymous):

@zzr0ck3r were you helping me or not?

OpenStudy (zzr0ck3r):

if you rationalize the denominator, its going to get ugly

OpenStudy (anonymous):

@zzr0ck3r well thats what the question asked me to do

OpenStudy (zzr0ck3r):

it also asked you to simplify, these are contradictions, ill show you

OpenStudy (anonymous):

ok thanks @zzr0ck3r

OpenStudy (zzr0ck3r):

\[\frac{\sqrt[3]{5a^5}}{\sqrt[3]{3b}}=\frac{\sqrt[3]{5a^5}}{\sqrt[3]{3b}}*\frac{\sqrt[3]{(3b^2)}}{\sqrt[3]{(3b)^2}}\] = \[\frac{\sqrt[3]{5a^5}*\sqrt[3]{(3b)^2}}{3b}=\frac{\sqrt[3]{5a^59b^2}}{3b}\] = \[\frac{\sqrt[3]{45a^3a^2b^2}}{3b}=a\frac{\sqrt[3]{45a^2b^2}}{3b}\]

OpenStudy (zzr0ck3r):

so we rationalized the denominator, but this looks way worse in my opinion.

OpenStudy (anonymous):

i have that but it says my answer is wrong @zzr0ck3r

OpenStudy (zzr0ck3r):

@mathstudent55 did I mess something up here?

OpenStudy (zzr0ck3r):

@joemath314159 do you see any problems?

OpenStudy (zzr0ck3r):

I have no forgotten you, just waiting for someone else to come take a look. there are so many different ways to write these that I want to make sure I did everything right before we start rearranging again. because this is not fun in latex and it takes me way too long

OpenStudy (zzr0ck3r):

bump the question when you can

OpenStudy (anonymous):

ok @zzr0ck3r no problem

OpenStudy (mathstudent55):

@zzr0ck3r I got the same you did.

OpenStudy (zzr0ck3r):

ty for looking

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

I don't know what to do because my HW won't accept the answer and I got the same answer as you guys @zzr0ck3r

OpenStudy (mathstudent55):

How do you enter the answer? Are you showing it's a cubic root in a way that is understandable?

OpenStudy (zzr0ck3r):

@salce08 ?